Logiweb dictionary of append

Up Help

0 0 append

1 2 append(∗, ∗)

The pyk compiler, version 0.grue.20050502 by Klaus Grue GRD-2005-06-02.UTC:19:54:02.237435 = MJD-53523.TAI:19:54:34.237435 = LGT-4624458874237435e-6